OpenStudy (anonymous):

You just purchased a cellular phone and are trying to determine to which cell phone company you will give your business. When you contacted the Talks-A-Lot Company, they were offering a monthly plan of $40 for 600 minutes and $0.35 for each minute exceeding the 600 minutes. In the Sunday paper you see an ad for the Chat-Away Company, which offers a monthly plan of $50 for 600 minutes and $0.10 for each minute exceeding the 600 minutes. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the equal cost?

OpenStudy (anonymous):

1.Set up a linear system consisting of two equations. Assume you will talk for a minimum of 600 minutes. The first equation would be for the Talks-A-Lot Company. The total cost, y, equals the base fee plus cost per minute times the number of minutes exceeding 600 minutes. The second equation would be set up just like the first, only you need to use the information for the Chat-Away Company. Solve the linear system using the substitution method. Please make sure to solve for both x and y. Show all work. Answer the questions, using complete sentences. How many minutes would you have to talk over and above the 600 minutes for the cost to be the same with both companies? What would be the cost when the minutes are the same? If you plan to talk for 1000 minutes, which company should you hire? Please show your total cost for both companies to prove your answer.

Directrix (directrix):

40 + .35(m-600) = 50 + .1(m-600) m = 640 (approx) So, 40 minutes above Equal cost is : 40 + .35(640-600) = $54